Hi. I need help with a word problem please. One plane flys 20mph faster than another plane. One planes goes 255 miles and the other 285miles. How fast was each plane flying? Thanks so much for your help! My email address is rmillsfamily@comcast.net
Answer by scott8148(6628)

t = d / r ___ the times are equal

255 / r = 285 / (r + 20)

"cross" multiplying ___ 255r + 5100 = 285r

5100 = 30r